# Distributionally Robust Acceleration Control Barrier Filter for Efficient UAV Obstacle Avoidance

### v1

@@ V-A Experimental Setup
Figure 3: Experimental lab setup. (Left) The green Crazyflie UAV must avoid two dynamic obstacles (marked in red): an approaching UAV and a thrown object. (Right) The motion-capturing system provides position data of the UAV and the obstacles. A video of our experiments is available for review.
@@
We have taken various Crazyflie limitations, including size, actuation bounds (acceleration, braking, and maneuvers), system latency, and clearance requirements, into consideration for our setup. We have adopted a similar metric for evaluation as used for our simulation study, and additionally measure the avoid time $t_{a}$ and the detection time $t_{d}$. The avoid time is the duration from the start of the avoid action to the end of hovering. The detection time is the duration from the first detection of the obstacle until the avoidance command is issued.
We use a Linux-based, CPU-only laptop as the base station, which was configured to capture both the obstacles and the UAV from OptiTrack. The position data was integrated into MATLAB via a Python bridge on the base station.Communication between the base station and UAV was established via the Crazyflie MATLAB-Python API at 100 Hz using full-state setpoints (position, velocity, acceleration, and attitude). The hardware setup uses the same DR-ACBF configuration as for the simulation study (see Section IV-A), except for $\|a_{e}\|_{\infty}\leq 0.8\,\mathrm{m/s^{2}}$, $R_{\mathrm{uav}}=0.10$ m and $R_{\mathrm{obs}}=0.10$ m. Furthermore, $d_{cl}$ is set to $0.05$ m and $t_{hd}$ to 0.2 s with the braking configuration: time to stop $t_{\mathrm{s}}$ within 0.7 s and braking intensity $k_{\mathrm{s}}=0.2$.
@@ V-B Experimental Results
We conduct two hardware experiments. In the first experiment, the UAV hovers at a fixed position, avoids an obstacle thrown by a human, and returns to the hovering position. In the second experiment, the hovering UAV avoids a directly approaching UAV and a human-thrown object.
TABLE IV: Crazyflie avoiding one dynamic obstacle with DR-ACBF.
Flight $v_{cl}$ $t_{r}$ $t_{cm}$ $t_{a}$ $t_{d}$ $d_{s}$
@@
6 2.96 66.1 53.2 249 2533 0.202
7 2.87 66.9 48.7 200 3616 0.241
8 3.04 66.4 64.8 200 3866 0.276
9 2.82 66.8 64.4 216 3683 0.337
10 3.11 66.4 59.1 250 3033 0.213
Mean 3.03 66.6 59.2 231 3277 0.366
Table IV presents the results of 10 flights of the first experiment using our DR-ACBF framework. In all flights, the UAV successfully avoided the thrown obstacle, which had a closing velocity $v_{cl}$ between 2.26 m/s and 3.66 m/s. The time measurements demonstrate the responsiveness and ability to execute avoidance actions efficiently. The reaction time $t_{r}$ was less than $67$ ms in 9 flights, only in flight #5 it was measured as 82.8 ms. Similarly, the average command time $t_{cm}$ was 59.18 ms, while the mean avoidance duration was 0.231 ms. In all flights, the detection times $t_{d}$ were above 2.5 s with a mean of 3.277 s, indicating a robust obstacle detection and tracking as well as an accurate response only when the obstacle is classified as unsafe. The separation distance $d_{s}$ between the UAV and the obstacle was always larger than 0.2 m with a mean of 0.365 m, ensuring collision-free operation within the tested environment. Notably, flights #1 and #4 exhibited the safe separation distances (0.205 m and 0.239 m, respectively), corresponding to the slowest and fastest velocities (2.26 m/s and 3.66 m/s).
@@
Table V presents the results of the second experiment, where the UAV successfully avoided the two obstacles in all five flights. For the approaching UAV (obstacle 1), the average closing velocity $v_{cl}$ was 2.35 m/s. We measured similar reaction and command times as in the first experiment (mean values for $t_{r}$ of 61.54 ms and for $t_{cm}$ of 41.50 ms). The avoidance duration was on average 0.18 s, while the mean detection time $t_{d}$ was 0.67 s. The mean separation distance $d_{s}$ was 0.325 m, confirming safe maneuvering performance. For obstacle 2, which involved a more unpredictable human-thrown object, the average $v_{cl}$ decreased to 1.49 m/s. The mean reaction and command times $t_{r}$ and $t_{cm}$ increased to 99.42 ms and 78.66 ms, respectively, while the average avoidance duration $t_{a}$ slightly increased to 0.217 s. The mean detection time $t_{d}$ significantly increased to 4.25 s, indicating that the object was detected early and entered the avoidance trigger after some time. Despite these changes, the UAV maintained a safe separation $d_{s}$ of 0.685 m from obstacle 2.
Finally, an important finding of the hardware experiments was that the separation distance $d_{s}$ was above the $R_{\mathrm{sum}}$ bound of 0.200 m. DR-ACBF was able to satisfy the specified safety constraint for all flights.
VI CONCLUSIONS
In this paper, we presented a distributionally robust acceleration-level safety filter that enforces collision avoidance through linear, half-space constraints on the commanded acceleration. Our DR-ACBF attains real-time avoidance without solving QPs. Simulation studies against three baseline methods demonstrate that DR-ACBF achieves comparable success rates and safety margins, while reducing computation and command times, thereby improving responsiveness for fast DOA. The sensitivity analysis confirms robustness to key design parameters, with high success rates. An ablation study further demonstrates that the DR-CVaR trigger is crucial for risk-aware operation. Crazyflie experiments demonstrate that DR-ACBF achieves reliable DOA on resource-limited platforms bridging the gap between formal safety and practical real-time performance. Future work will extend this framework to integrate RADAR or vision sensing modality with a medium-sized UAV platform to assess the applicability of the proposed framework.

### v2

@@ V-B Experimental Results
TABLE IV: Crazyflie avoiding one dynamic obstacle with DR-ACBF in UAV hovering and moving forward condition.
Condition Flight $v_{cl}$ $t_{r}$ $t_{cm}$ $t_{a}$ $t_{d}$ $d_{s}$
@@
Mean 3.10 66.6 60.3 239 3209 0.479
Moving 1 3.50 84.2 74.5 218 2599 0.312
2 3.74 84.0 71.6 241 2725 0.258
3 3.49 67.2 52.3 208 2616 0.446
4 3.36 82.9 70.3 225 2923 0.414
5 2.90 75.6 63.1 226 2678 0.324
Mean 3.40 78.8 66.4 223 2708 0.351
We conduct three hardware experiments. In the first two experiments, the UAV hovers at a fixed position and moves toward a waypoint, respectively. In both experiments, it avoids a single obstacle thrown by a human. In the third experiment, the hovering UAV avoids two obstacles: a directly approaching UAV and a human-thrown object. Table IV presents the results of 5 flights of the first and second experiment using our DR-ACBF framework. In all flights, the UAV successfully avoided the thrown obstacle, which had a closing velocity $v_{cl}$ between 2.26 m/s and 3.74 m/s. The reaction time $t_{r}$ was shorter in the hovering ($66.6$ ms) than in the moving ($78.8$ ms) experiment. The mean command time $t_{cm}$ showed a similar trend ($60.3$ ms vs. $66.4$ ms). Whereas the mean avoidance duration $t_{a}$ and detection times $t_{d}$ were smaller for the moving experiment ($223$ ms and $2.7$ s vs. $239$ ms and $3.2$ s). This suggests a reliable trigger, as the intervention occurs only when the obstacle becomes unsafe. The separation distance $d_{s}$ between the UAV and the obstacle was always greater than the specified $R_{sum}=0.2$ m, with a mean of $0.479$ m for hovering and $0.351$ m moving experiment, ensuring collision-free operation within the tested environment. The simulation $t_{cm}$ is smaller than the hardware due to conservative safety settings. Hardware presents additional delays due to communication and actuation limitations imposed by low battery levels.
@@
Table V presents the results of the third experiment, where the UAV successfully avoided the two obstacles in all five flights. For the approaching UAV (obstacle 1), the average closing velocity $v_{cl}$ was 2.35 m/s. We measured similar reaction and command times as in the first experiment (mean values for $t_{r}$ of 61.5 ms and for $t_{cm}$ of 41.5 ms). The avoidance duration was on average 0.18 s, while the mean detection time $t_{d}$ was 0.67 s. The mean separation distance $d_{s}$ was 0.325 m, confirming safe maneuvering performance. For obstacle 2, which involved a more unpredictable human-thrown object, the average $v_{cl}$ decreased to 1.49 m/s. The mean reaction and command times $t_{r}$ and $t_{cm}$ increased to 99.42 ms and 78.66 ms, respectively, while the average avoidance duration $t_{a}$ slightly increased to 0.217 s. The mean detection time $t_{d}$ significantly increased to 4.25 s, indicating that the object was detected early and entered the avoidance trigger after some time. Despite these changes, the UAV maintained a safe separation $d_{s}$ of 0.685 m from obstacle 2. Finally, an important finding of the hardware experiments was that the separation distance $d_{s}$ was above the $R_{\mathrm{sum}}$ bound of 0.2 m. DR-ACBF was able to satisfy the specified safety constraint for all flights.
VI Conclusion
In this paper, we presented a distributionally robust acceleration-level safety filter that enforces collision avoidance through linear, half-space constraints on the commanded acceleration. Our DR-ACBF attains real-time avoidance without solving QPs. Simulation studies against three baseline methods demonstrate that DR-ACBF achieves comparable success rates and safety margins, while reducing computation and command times, thereby improving responsiveness for fast DOA. The sensitivity analysis confirms robustness to key design parameters, with high success rates. An ablation study further demonstrates that the DR-CVaR trigger is crucial for risk-aware operation, and projection is much lighter than QP-based ACBF. Crazyflie experiments demonstrate that DR-ACBF achieves reliable DOA on resource-limited platforms, bridging the gap between formal safety and practical real-time performance. Future work will extend this framework to integrate RADAR or vision sensing modality with a medium-sized UAV to assess the applicability of the proposed framework.